OpenStudy (anonymous):

Find the slope of the line that passes through (3, -4) and (-1, 2). A. 2/3 B. 3/2 C. -2/3 D. -3/2

OpenStudy (anonymous):

slope = y1 - y2/x1 - x2 = -4-2/3-(-1) = -6/4 = -3/2

OpenStudy (anonymous):

use the slope formula which is (x2-x1)/(y2-y1)

OpenStudy (anonymous):

x1=3 x2=-1 y1=-1 y2=2

OpenStudy (anonymous):

^^^^noo it's delta y/delta x